FIND THE CHANCE OF THROWING MORE THAN 15 IN ONE THROW WITH 3 DICE.
Answer by Fombitz(32388)

Throwing 3 dice gives a possible 216 outcomes.
Of those, 10 sum to greater than 15.
4 6 6
5 5 6
5 6 5
5 6 6
6 4 6
6 5 5
6 5 6
6 6 4
6 6 5
6 6 6
So the probability,
P(Sum>15)=10/216=5/108